# A Fair and Lightweight Consensus Algorithm for IoT

### v1

@@ 3.1 Experiment objectives and setting
Our prime objective is to evaluate the performance of the proposed consensus algorithm concerning its fairness property and its robustness to attackers. To this end, we have developed a reference implementation of the algorithm in Python. To enable the communication and message exchange between the nodes, we implemented a REST API using Flask. Over this REST API a simple gossip protocol was developed, allowing nodes to spread messages to five random neighbors. Avoidance of repeating already sent messages, such as block proposals, was implemented to reduce the communication overhead. It is important to note that in terms of performance Python and Flash are sub-optimal choices. However, their ease of use enables rapid development which fits our objective of creating a Proof-of-Concept (PoC) to assess the algorithm’s fundamental properties.
Regarding the configuration of the consensus algorithm, we made the following choices. Similarly to other algorithms [28, 18], We set the overall round timeout period to 18 seconds (12 seconds for the block proposal, and 8 seconds for voting). Empirically, this time window would be sufficient for spreading information (i.e., messages) over the network. Furthermore, to reduce the network overhead we also set $THRESHOLD_{proposal}=0.9$, allowing only 10% of the nodes to propose a block in each round. Regarding the voting mechanism, we set $THRESHOLD_{reputation}=0.8$ (in the range [0,1]) to allow only high-reputable nodes to vote, while $Threshold_{vote}$ is set to allow a voting committee of 11 nodes. Nodes obtain reputation information via a reputation system. We also note that we set the size of blocks to 200 kilobytes. Lastly, to reduce the computation overhead introduced by cryptographic operations (e.g., digital signatures of messages), we replaced such operations with less computationally intensive alternatives (e.g., time delays).
For the experiment, we deployed our sample implementation in the AWS cloud infrastructure, deploying up to 100 nodes (depending on the experiment use-case) in a similar number of t3.micro instances. Out of these nodes, 15 are highly reputable nodes, with a reputation of 0.9, from which the voting committee members will be randomly selected. Note that despite that the number of the highly-reputable nodes does not have an impact on the experiment results, it should be high enough to allow the formation of a voting committee. We assume that these nodes remain honest throughout the experiments. We acknowledge that our experiment setting, and implementation are not optimized for computational and network efficiency. However, we argue that this has a limited impact on our evaluation since we study the behavior of our algorithm with respect to fairness and robustness to adversaries rather than for efficiency and throughput.
@@ 3.2 Assessing the Fairness property
The first finding of our experiment is that the algorithm maintains a stable throughput of three blocks per minute, regardless of the number of nodes. Although our setup involves a relatively small scale (up to 100 nodes) compared to large public blockchain networks with thousands of nodes, the results indicate that the proposed approach performs as expected, with a round time of approximately 20 seconds. Additionally, most nodes stay synchronized with the global chain, though a slight drop in synchronization is observed as the network size increases.After inspecting the logs of the nodes, we found that the desynchronized nodes were lagging for one block, which could be a result of network discrepancies. These nodes are expected to catch up with the last block during the next round. Lastly, regarding the fairness property of our consensus algorithm, we notice that the block proposer diversity grows from 76% to 90% as the number of nodes increases. This behavior is expected as the uniform distribution of the probability of the lotteries is better expressed on a greater scale.
@@ 3.3 Robustness to attacks
Before we move on to discuss how we assess the robustness of attacks, we need to define the attackers’ profile. To this end, we consider that an attacker can deploy multiple malicious nodes that will not respect the consensus protocol. Since attackers cannot influence the integrity of the exchanged messages, and cannot forge the random lots (due to the use of VRFs), they concentrate their efforts on selecting which information to disseminate over the network to affect the block added to the chain. Malicious nodes may selectively forward block proposals and votes aiming at stopping the expansion of the chain with new blocks. Next to that, malicious nodes aim at increasing the chances that a malicious node is chosen to propose a block, effectively halting the addition of a block for that round. In our experiments, we assume that malicious nodes cannot have a high reputation and thus cannot participate in the voting committee.
To assess the robustness of our algorithm against attackers, we employ similar metrics to the ones used in the previous section. In particular, we measure how the presence of attackers impacts the fairness and the expansion of the chain. We do this with a set of experiments deploying 100 nodes and a varying ratio of malicious nodes on the network. For each case, we let the consensus algorithm run for ten minutes. We note that by definition, the malicious nodes are deliberately not in sync with the chain. The results of our experiments are shown in Table 2
@@
Evaluating the results, we observe that the ratio of malicious nodes has an impact on the throughput of the blockchain network. In particular, we notice a small drop of 14% in the number of blocks added when the ratio of attackers is relatively low (10% to 30%), however, a more significant drop of 40% -50% is noticed when the number of malicious nodes represents almost half of the nodes. This behavior can be justified by the fact that the ratio of malicious nodes in the network is proportional to the probability of being selected to propose a block. Such an effect could be mitigated by using the reputation system to decrease the reputation of nodes that consistently misbehave, and eventually ban them from the network. We leave the investigation of this use case for future work. We note that regardless of the drop in overall throughput, the network manages to operate with both safety and finality. Last but not least, it is noticed that the block proposer diversity (i.e., fairness) and the (honest) nodes synchronized to the global chain are not affected by the presence of adversaries in the network. This is evident in the scenario with 50 malicious nodes, where 14 out of 15 blocks were proposed by different nodes, and 47 out of 50 nodes remained synchronized.
4 Related work
Consensus algorithms are regarded as the cornerstone of blockchain as they provide the means for nodes to agree on the state of the blockchain. Specifically, they provide the mechanisms based on which nodes agree on the blocks to be added to the blockchain. The characteristics of the consensus algorithm are crucial both for the performance and security of properties on the blockchain solution. To this end, several different approaches have been proposed in the literature originating from the domain of distributed network systems and recently in blockchain technology per se. In general, consensus algorithms can be classified into two broad categories: voting-based and proof-based [19].
Traditional distributed systems approaches have been adapted for permissioned blockchains, as they share similar foundational assumptions. These methods rely on a voting mechanism where nodes elect a leader and vote on block proposals, with final decisions determined by majority rule. Their primary goal is to ensure consistency and finality, guaranteeing that all nodes have a unified view of the blockchain state after each round. However, these approaches tend to be more centralized and face scalability issues due to the high communication overhead required for consensus. Additionally, many voting-based models assume that most participating nodes are both honest and consistently available.
Over the past decades, Paxos [22] was the dominant voting-based consensus mechanism, although its complexity hindered its application. To address this issue, Raft [30] was proposed as a simple and understandable approach to reaching consensus. Raft addressed the issue of understandability by introducing a straightforward leader election mechanism, in which each node has the chance to campaign for a Leader role for a particular round. This was achieved by having random time-based rounds, where a node could request the support (i.e., votes) of its peers, aiming for a majority. Both approaches although being crash-fault tolerant assumed the honesty (i.e., trust) of all the participating nodes, restricting their application in many real-case scenarios. To this end, Byzantine-fault tolerant (BFT) solutions were proposed. One of the most widely applied and efficient mechanisms of this kind is Practical Byzantine Fault Tolerance (pBFT) [9] consensus algorithm which was designed to tolerate up to |(n-1)/3| malicious nodes. It uses a three-phase protocol, where the failure to reach consensus among all consensus nodes triggers the election of a new Primary node (leader). The major drawback of pBFT is its quadratic complexity $O(n^{2})$ compared to the linear complexity of the other two algorithms. Modern and scalable voting-based approaches have been also proposed exclusively for blockchain networks. A prominent example is Proof-Of-Vote [24] in which a committee of predefined members elects butler nodes, which are responsible for validating transactions and adding blocks to the chain.
Regarding public blockchains, Proof of Work (PoW) [15, 8] is the most popular consensus algorithm, since it is the one adopted by Bitcoin. It is a fully decentralized consensus mechanism that requires the network members to put computational effort into solving an arbitrary mathematical puzzle. The first node to solve the puzzle adds a block to the chain. Confirming a solution to such a puzzle can be easily and undeniably verified by other nodes. PoW approaches, however, suffer from low throughput and have a significant environmental impact due to the high energy consumption required. To address the latter, different consensus mechanisms have been proposed in the literature, such as Proof of Capacity / Proof of Storage [4, 16] that pose a challenge related to the storage of data, which has a lower environmental footprint. Other approaches focus on selecting nodes in a physical manner employing the notion of randomness similar to participating in a lottery. For instance, in Proof of Elapsed time [12] each node has a random timer and the node that manages to time out first adds a block to the chain. Similarly, in the Proof of Luck (PoL) algorithm [28], each node calculated a random number and the luckiest (e.g., smallest number) adds a block. To guarantee the honest behavior of the nodes and the indisputably of the random procedures, a Trusted execution environment (TEE) is required to be present in each node. Based on the previous discussion it becomes apparent that these solutions require nodes to have a considerable number of resources or features, which renders them inappropriate for an IoT context [41, 38, 35, 33]. Nonetheless, recent approaches like Proof of Verifiable Function (PoVF) utilize Verifiable Delay Functions (VDFs) and VRFs to achieve unpredictable node selection without relying on specialized hardware [42].
Apart from approaches that rely on a physical challenge to select a node for adding a block to the chain, other consensus algorithms make such a choice based on node properties. For instance, Proof of Authority [37] solutions rely on a set of pre-authorized nodes, whose identity and trust have already been verified. In each round, a node from that set is selected to add a block, while the rest of the nodes in the network follow that decision. Although quite efficient, these approaches increase the level of centralization and authorized nodes usually must be pre-approved in the physical world [41, 38]. The most prominent consensus algorithm in the aforementioned family is Proof of Stake (PoS) [21], which is quickly gaining ground on modern crypto coins. In this approach, the node to add a block to the chain is selected considering the stake (e.g., the coins held) that a node has in the network. The underlying intuition is straightforward, the more the stake, the greater the chances to be selecting the node responsible for adding a block. In this context, nodes with high stakes have an incentive for the network to operate as designed and be profitable. Moreover, nodes who do not respect their responsibilities may lose part of their stake (i.e., slashing) as punishment, thus they will not misbehave. PoS algorithms exist in many different flavors, such as Delegated PoS [3] which aim to be more efficient or Pure PoS [18, 13] solutions which aim to be more secure and fair. The main drawback of such approaches is that they are designed for cryptocurrencies where nodes have an economic incentive, making their application difficult in different contexts [41, 38]. Moreover, such algorithms favor nodes with significant investments in the network, leading to increased centralization.
Although narrow in their applicability, PoS algorithms were proven to be secure, efficient, and with limited resource requirements, thus they gained the interest of the research community. To address their applicability issues, researchers have worked on expanding their concept to cover a wider variety of properties used in the consensus. To this end, Proof of Importance (PoI) [29] algorithms have been proposed. This family of algorithms extends the concept of stake by considering other aspects that highlight the importance of a node in the network. For example, the contribution to the network, the number of successful transactions, and the trustworthiness of the nodes are considered. A particular type of PoI type of algorithm is Proof of Trust / Reputation [44, 20, 36] algorithms which assess the overall trustworthiness of nodes (e.g., based on their characteristics and past behavior) to select a node for adding a block to the chain. These approaches are quite promising and there is considerable ongoing research in the field of IoT. However, the main challenge regarding these algorithms is related to their fairness, as nodes with higher importance tend to dominate the blockchain.
In real-world applications, most consensus algorithms adopt a hybrid approach, combining multiple mechanisms to capitalize on their strengths while mitigating their shortcomings. One such example is Proof of Activity [6], which integrates elements of both PoW and PoS. In this model, a lightweight PoW process is used to mine a block, after which PoS is used for validation and block addition. Other notable hybrid approaches include Ethereum’s Proof-of-Stake and Algorand [18, 13], which, while primarily PoS-based, incorporate pBFT voting schemes to enhance security and ensure stronger finality guarantees. Hybrid consensus mechanisms have also been explored in the IoT domain, such as PoEWAL [2], which combines PoW with PoLuck. Given their ability to balance security, efficiency, and scalability, hybrid approaches present a promising direction for IoT applications.
While we have already discussed the applicability of fundamental consensus paradigms in IoT, several recent studies have developed algorithms exclusively for IoT [33]. For example, the authors in [7] extend the HyperLedger Fabric framework by proposing a novel approach to reduce the number of participating nodes in the consensus process, enhancing its suitability for IoT environments. In [14], a leader-based consensus algorithm is introduced, where leaders precompute a hash-based token (i.e., consensus code) and are assigned to manage transactions associated with that token. This approach allows validators to create and commit blocks contraining transactions under their supervision, reducing the need for extensive coordination among nodes.
Additionally, two works share similarities with our proposed algorithm: REVO [5] and LVCA [40]. REVO leverages reputation to form trusted committees that elect nodes for block addition, whereas LVCA employs a random lottery voting-based consensus among trusted nodes to select the block proposer. In contrast, our algorithm emphasizes fairness by randomly selecting block proposers from the global pool of nodes while using reputation solely for validation and block addition to the chain.
@@ 5 Conclusion & Future Work
This work introduced a fair and lightweight hybrid consensus algorithm designed for IoT environments, balancing security, fairness, and efficiency. It combines random lotteries for fairness with reputation-based voting for safety and finality, all with minimal resource demands. A sample implementation validated the algorithm’s fundamental fairness property and its resilience against adversarial attacks. However, real-world efficiency and scalability in the field of IoT remain open for further investigation. In addition, future research could further leverage reputation mechanisms employed to enhance security and throughput without compromising fairness.

### v2

@@ 4.1 Experiment objectives and setting
Our primary objective is to evaluate the performance of the proposed consensus algorithm concerning its fairness property and its robustness to attackers. To this end, we have developed a reference implementation of the algorithm in Python. To enable communication and message exchange between the nodes, we implemented a REST API using Flask. Over this API a simple gossip protocol was developed, allowing nodes to spread messages to five random neighbors. Avoidance of repeating already sent messages, such as block proposals, was implemented to reduce communication overhead. It is important to note that in terms of performance, Python and Flash are sub-optimal choices. However, their ease of use enables rapid development which fits our objective of creating a Proof of Concept (PoC) to assess the algorithm’s fundamental properties.
Regarding the configuration of the consensus algorithm, we made the following choices. Similarly to other algorithms [milutinovic2016proof, gilad2017algorand], we set the overall round timeout period to 20 seconds (12 seconds for the block proposal and 8 seconds for voting). Empirically, this time window would be sufficient for spreading information (i.e., messages) over the network. Furthermore, to reduce the network overhead, we also set $THRESHOLD_{proposal}=0.9$, allowing only 10% of the nodes to propose a block in each round. Regarding the voting mechanism, we set $THRESHOLD_{reputation}=0.8$ (in the range [0,1]) to allow only high-reputable nodes to vote, while $Threshold_{vote}$ is set to allow a voting committee of 11 nodes. The nodes obtain reputation information via a reputation system. We also note that we set the size of the blocks at 200 kilobytes. Lastly, to reduce the computation overhead introduced by cryptographic operations (e.g., digital signatures of messages), we replaced such operations with less computationally intensive alternatives (e.g., time delays).
For the experiment, we deployed our sample implementation in the AWS cloud infrastructure, deploying up to 100 nodes (depending on the experiment use-case) in a similar number of t3.micro instances. Of these nodes, 15 are highly reputable, with a reputation of 0.9, from which the voting committee members will be randomly selected. Note that despite that the number of highly reputable nodes does not have an impact on the experiment results, it should be high enough to allow the formation of a voting committee. We assume that these nodes remain honest throughout the experiments. We also acknowledge that our experiment setting and implementation are not optimized for computational and network efficiency. However, we argue that this has a limited impact on our evaluation since we study the behavior of our algorithm with respect to fairness and robustness to adversaries rather than for efficiency and throughput.
@@ 4.2 Assessing the Fairness property
The first finding of our experiment is that the algorithm maintains a stable throughput of three blocks per minute, regardless of the number of nodes. Despite the relatively small-scale setup (up to 100 nodes) compared to large public blockchain networks with thousands of participants, the results indicate that the proposed approach performs as expected, with a round-time of approximately 20 seconds. Furthermore, most nodes stay synchronized with the global chain, though a slight drop in synchronization is observed as the network size increases. After inspecting the node logs, we found that the desynchronized nodes were lagging for a block, which could be the result of network discrepancies. These nodes are expected to catch up with the last block during the next round. Lastly, regarding the fairness property of our consensus algorithm, we notice that the block proposer diversity grows from 76% to 90% as the number of nodes increases. This behavior is expected because the uniform distribution of the probability of the lotteries is better expressed on a greater scale.
@@ 4.3 Robustness to attacks
Before we move on to discuss how we assess the robustness of attacks, we need to define the attacker’s profile. In this regard, we consider that an attacker can deploy multiple malicious nodes that will not respect the consensus protocol. Since attackers cannot influence the integrity of the exchanged messages and cannot forge the random lots (due to the use of VRFs), they concentrate their efforts on selecting which information to disseminate over the network to affect the block added to the chain. Malicious nodes may selectively forward block proposals and votes aiming to stop the expansion of the chain with new blocks. Next to that, malicious nodes aim to increase the chances that a malicious node is chosen to propose a block, effectively halting the addition of a block for that round. In our experiments, we assume that malicious nodes cannot have a high reputation and thus cannot participate in the voting committee.
To assess the robustness of our algorithm against attackers, we employ metrics similar to those used in the previous section. In particular, we measure how the presence of attackers impacts the fairness and expansion of the chain. We do this with a set of experiments deploying 100 nodes and a varying ratio of malicious nodes on the network. For each case, we let the consensus algorithm run for ten minutes. We note that by definition, the malicious nodes are deliberately not in sync with the chain. The results of our experiments are shown in Table 2
@@
Evaluating the results, we observe that the ratio of malicious nodes has an impact on the throughput of the blockchain network. In particular, we notice a small drop of 14% in the number of blocks added when the ratio of attackers is relatively low (10% to 30%), however, a more significant drop of 40% -50% is noticed when the number of malicious nodes represents almost half of the nodes. This behavior can be justified by the fact that the ratio of malicious nodes in the network is proportional to the probability of being chosen to propose a block. Such an effect could be mitigated by using the reputation system to decrease the reputation of nodes that consistently misbehave, and eventually ban them from the network. We leave the investigation of this use case for future work. We note that regardless of the drop in overall throughput, the network manages to operate with both safety and finality. Last but not least, it is noticed that the block proposer diversity (i.e., fairness) and the (honest) nodes synchronized to the global chain are not affected by the presence of adversaries in the network. This is evident in the scenario with 50 malicious nodes, where 14 of 15 blocks were proposed by different nodes, and 47 out of 50 nodes remained synchronized.
@@ 5 Conclusion & Future Work
This work introduced a fair and lightweight hybrid consensus algorithm designed for IoT environments, balancing security, fairness, and efficiency. It combines random lotteries for fairness with reputation-based voting for safety and finality, all with minimal resource demands. A sample implementation validated the algorithm’s fairness property and its resilience against adversarial attacks. However, real-world efficiency and scalability in IoT remain open for further investigation. Finally, future research could further leverage the reputation mechanism to improve security and performance without compromising fairness.
Acknowledgment
This work has received funding from the EU Horizon Europe Programme in the framework of the "Upscaling Innovative Green Urban Logistics Solutions Through MultiActor Collaboration and PI-Inspired Last Mile Deliveries" project (URBANE), under GA Number 101069782